# Optimal Transport for $\epsilon$-Contaminated Credal Sets

### v1

@@ 3.1 Extension
Suppose $\mathcal{X}$ and $\mathcal{Y}$ are separable, so that the elements of $\Delta_{\mathcal{X}}$ and $\Delta_{\mathcal{Y}}$ are Radon measures. Let $c:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}_{+}$ be a Borel-measurable (cost) function. Given lower probabilities $\underline{P}$ and $\underline{Q}$ on $\mathcal{X}$ and $\mathcal{Y}$, respectively, we want to find the joint lower probability $\underline{\alpha}$ (also called the lower optimal transport plan) on $(\mathcal{X}\times\mathcal{Y},\mathcal{F}\times\mathcal{G})$ that attains the infimum
$\inf\left\{{\int_{\mathcal{X}\times\mathcal{Y}}c(x,y)\underline{\alpha}(\text{d}(x,y)):\underline{\alpha}\in\Gamma(\underline{P},\underline{Q})}\right\},$ (11)
@@
In Definition 9, we assume that function $(x,y)\mapsto c(x,y)$ is Choquet integrable with respect to $\underline{\alpha}$, for all $\alpha\in\Gamma(\underline{P},\underline{Q})$ (Definition 3, Proposition 4). In Imprecise Probability theory (Augustin et al., 2014; Troffaes and de Cooman, 2014; Walley, 1991) there is not a unique way to perform conditioning (Gong and Meng, 2021; Caprio and Seidenfeld, 2023; Caprio and Gong, 2023), so we need to be extra careful when defining $\Gamma(\underline{P},\underline{Q})$ in (11). In this work, we consider the joint lower probabilities resulting from geometric conditioning, and write $\Gamma(\underline{P},\underline{Q})\equiv\Gamma^{\text{geom}}(\underline{P},\underline{Q})$. In that case, the conditional lower probability resulting from a joint probability $\underline{G}$ is derived as
$\underline{G}(A\mid B)=\frac{\underline{G}(A,B)}{\underline{G}_{\mathcal{Y}}(B)},\quad\forall A\in\mathcal{F},\forall B\in\mathcal{G},$ (12)
where we call $\underline{G}_{\mathcal{Y}}$ the marginal lower probability of $G$ on $\mathcal{Y}$, and similarly for $\underline{G}_{\mathcal{X}}$. In the remainder of this paper, we will omit the subscript when it is clear from the context which marginal of joint lower probability $\underline{G}$ we are considering. If the marginal lower probabilities correspond to the lower envelopes of $\epsilon$-contaminated credal sets, then using joint lower probabilities that can be decomposed as in (12) entails that the elements of $\Gamma(\underline{P},\underline{Q})$ are such that, for all $A\in\mathcal{F}$ and all $B\in\mathcal{G}$,
@@
and
$\underline{G}(B\mid A)=\frac{\underline{G}(B,A)}{\underline{P}(A)}=\frac{\underline{G}(A,B)}{(1-\epsilon)P(A)}.$
We can then consider a restricted version of LPK. [^6] In Definition 10, we assume that function $(x,y)\mapsto c(x,y)$ is Choquet integrable with respect to $\underline{\alpha}$, for all $\alpha\in\Gamma_{R}(\underline{P},\underline{Q})$
@@
Suppose $\mathcal{X}$ and $\mathcal{Y}$ are separable, so that the elements of $\Delta_{\mathcal{X}}$ and $\Delta_{\mathcal{Y}}$ are Radon measures. Let $c:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}_{+}$ be a Borel-measurable (cost) function. Given lower probabilities $\underline{P}$ and $\underline{Q}$ on $\mathcal{X}$ and $\mathcal{Y}$, respectively, we want to find the joint lower probability $\underline{\alpha}$ on $(\mathcal{X}\times\mathcal{Y},\mathcal{F}\times\mathcal{G})$ that attains the infimum
$\inf\left\{{\int_{\mathcal{X}\times\mathcal{Y}}c(x,y)\underline{\alpha}(\text{d}(x,y)):\underline{\alpha}\in\Gamma_{R}(\underline{P},\underline{Q})}\right\},$ (13)
where $\Gamma_{R}(\underline{P},\underline{Q})\subset\Gamma(\underline{P},\underline{Q})$ is the collection of all joint lower probabilities (i) that can be written as an $\epsilon$-contamination of countably additive joint probability measures $G\in\Delta_{\mathcal{X}\times\mathcal{Y}}$, and (ii) whose marginals on $\mathcal{X}$ and $\mathcal{Y}$ are $\underline{P}$ and $\underline{Q}$, respectively.
@@
If $\underline{P}$ and $\underline{Q}$ are the lower envelopes of the $\epsilon$-contaminations $\mathcal{P}_{\epsilon}\subseteq\Delta_{\mathcal{X}}$ and $\mathcal{Q}_{\epsilon}\subseteq\Delta_{\mathcal{Y}}$ of $P\in\Delta_{\mathcal{X}}$ and $Q\in\Delta_{\mathcal{Y}}$, respectively, then the RLPK of Definition 10 is equivalent to the classical Kantorovich’s OT Problem involving $P$ and $Q$.
@@
$\underline{G}(A,B)$ $=\underline{G}(A\mid B)\underline{G}(B)=\frac{G(A,B)}{Q(B)}(1-\epsilon)Q(B)$
$=\underline{G}(B\mid A)\underline{P}(A)=\frac{G(A,B)}{P(A)}(1-\epsilon)P(A)=(1-\epsilon)G(A,B).$
@@
So we can write $\Gamma_{R}(\underline{P},\underline{Q})=(1-\epsilon)\Gamma(P,Q)=\{(1-\epsilon)G:G\in\Gamma(P,Q)\}$, where set $\Gamma(P,Q)$ is the the collection of all (countably additive) probability measures on $\mathcal{X}\times\mathcal{Y}$ whose marginals on $\mathcal{X}$ and $\mathcal{Y}$ are ${P}$ and ${Q}$, respectively. This shows that $\Gamma_{R}(\underline{P},\underline{Q})$ is nonempty if and only if $\Gamma(P,Q)\neq\emptyset$. In turn,
@@
$\inf\left\{{\int_{\mathcal{X}\times\mathcal{Y}}c(x,y)\underline{\alpha}(\text{d}(x,y)):\underline{\alpha}\in\Gamma_{R}(\underline{P},\underline{Q})}\right\}$
@@
$=\inf\left\{{(1-\epsilon)\int_{\mathcal{X}\times\mathcal{Y}}c(x,y){\alpha}(\text{d}(x,y)):{\alpha}\in\Gamma(P,Q)}\right\}$ (14)
@@
$=\inf\left\{{\int_{\mathcal{X}\times\mathcal{Y}}c(x,y){\alpha}(\text{d}(x,y)):{\alpha}\in\Gamma(P,Q)}\right\},$ (15)
@@
where (14) comes from Proposition 4 and our definition of $\Gamma_{R}(\underline{P},\underline{Q})$, and the last equality comes from $c$ being positive. The fact that the optimization problem in (15) is the classical Kantorovich’s OT Problem (Kantorovich, 1942) concludes our proof. ∎
Similarly to Theorem 8, Theorem 11 tells us that RLPK is equivalent to a slightly more restrictive version of the classical Kantorovich’s OT problem, since we need to require that cost function $c$ is Choquet integrable with respect to the elements of $\Gamma_{R}(\underline{P},\underline{Q})$. [^7] If $c$ is bounded, we know that this is the case by Proposition 4. As pointed out earlier, though, for all practical purposes this does not entail loss of generality.
@@
We now give sufficient conditions for the minimizer of (13) to exist, in the context of $\epsilon$-contaminated credal sets.
@@
Let $\underline{P}$ and $\underline{Q}$ be the lower probabilities of Theorem 11. Endow $\mathcal{X}\times\mathcal{Y}$ with a topology $\tau$ that is contained in ×$\mathcal{F}\times\mathcal{G}$. That is, every open subset of $\mathcal{X}\times\mathcal{Y}$ is a measurable set, and ×$\mathcal{F}\times\mathcal{G}$ is at least as fine as the Borel $\sigma$-algebra on $\mathcal{X}\times\mathcal{Y}$. If $\Gamma_{R}(\underline{P},\underline{Q})$ is tight (with respect to topology $\tau$), and if cost function $c$ in (13) is also lower semicontinuous, then a minimizer for (13) exists.
@@
Ambrosio et al. (2005) show that if $\Gamma(P,Q)$ is $\tau$-tight, and if $c$ is lower semicontinuous, then there is a minimizer for the classical Kantorovich’s OT problem. Since $\Gamma_{R}(\underline{P},\underline{Q})=(1-\epsilon)\Gamma(P,Q)$, we know that if $\Gamma_{R}(\underline{P},\underline{Q})$ is $\tau$-tight, then so is $\Gamma(P,Q)$. The proof follows by the equivalence established in Theorem 11. ∎
@@
Let $\underline{P}$ and $\underline{Q}$ be the lower envelopes of the $\epsilon$-contaminations $\mathcal{P}_{\epsilon}\subseteq\Delta_{\mathcal{X}}$ and $\mathcal{Q}_{\epsilon}\subseteq\Delta_{\mathcal{Y}}$ of $P\in\Delta_{\mathcal{X}}$ and $Q\in\Delta_{\mathcal{Y}}$, respectively. When the minimizer $\underline{\alpha}$ of RLPK can be written in the form $\underline{\alpha}(\text{d}(x,y))=\underline{P}(\text{d}x)\delta_{y=T(x)}$, then $T$ is an optimal transport map and RLPK is equivalent to LPM.
@@
Given the way we defined $\underline{P}$ and $\underline{Q}$, we have that
$\underline{\alpha}(\text{d}(x,y))=\underline{P}(\text{d}x)\delta_{y=T(x)}$ $\iff(1-\epsilon){\alpha}(\text{d}(x,y))=(1-\epsilon){P}(\text{d}x)\delta_{y=T(x)}$
$\iff{\alpha}(\text{d}(x,y))={P}(\text{d}x)\delta_{y=T(x)}.$
Thorpe (2018, Section 1.2) shows that when the minimizer ${\alpha}$ of Kantorovich’s classical OT problem can be written in the form ${\alpha}(\text{d}(x,y))={P}(\text{d}x)\delta_{y=T(x)}$, [^8] Conditions sufficient for such a condition can be found in Thorpe (2018, Chapter 4). then $T$ is Monge’s OT map, and Monge’s and Kantorovich’s problems are equivalent. The proof, then, follows by Theorems 8 and 11. ∎
@@
Let $\underline{P}$ and $\underline{Q}$ be the lower envelopes of the $\epsilon$-contaminations $\mathcal{P}_{\epsilon},\mathcal{Q}_{\epsilon}\subseteq\Delta_{\mathbb{R}^{d}}$ of $P=\mathcal{N}_{d}(0,\Sigma_{P})$ and $Q=\mathcal{N}_{d}(0,\Sigma_{Q})$, two multivariate Normals on $\mathcal{X}=\mathcal{Y}=\mathbb{R}^{d}$, respectively. Select $c(x,y)=|y-Ax|^{2}/2$, where $A\in\mathbb{R}^{d\times d}$ is invertible. Then, the optimal map that solves LPM is
$x\mapsto T(x)=(A^{\top})^{-1}\Sigma_{P}^{-1/2}\left(\Sigma_{P}^{1/2}A^{\top}\Sigma_{Q}A\Sigma_{P}^{1/2}\right)^{1/2}\Sigma_{P}^{-1/2}x$ (16)
and the optimal plan that solves RLPK is $\underline{\alpha}(\text{d}(x,y))=\underline{P}(\text{d}x)\delta_{y=T(x)}$.
@@
Notice that $\mathbb{R}^{d}$ is separable. Galichon (2016) shows that if $P=\mathcal{N}_{d}(0,\Sigma_{P})$, $Q=\mathcal{N}_{d}(0,\Sigma_{Q})$, and $c(x,y)=|y-Ax|^{2}/2$, then Monge’s (classical) OT map is the one in (16), and also that
${\alpha}(\text{d}(x,y))={P}(\text{d}x)\delta_{y=T(x)},$ (17)
so that Monge’s and Kantorovich’s (classical) problems are equivalent.
Now, by Theorem 8, we have that if $\underline{P}$ and $\underline{Q}$ are lower envelopes of the $\epsilon$-contaminations $\mathcal{P}_{\epsilon},\mathcal{Q}_{\epsilon}\subseteq\Delta_{\mathbb{R}^{d}}$ of $P=\mathcal{N}_{d}(0,\Sigma_{P})$ and $Q=\mathcal{N}_{d}(0,\Sigma_{Q})$, respectively, then LPM coincides with the classical Monge’s OT Problem involving $P$ and $Q$. In turn, this implies that the map in (16) is also the OT map between $\underline{P}$ and $\underline{Q}$. In addition, by Theorem 11, we know that RLPK coincides with the classical Kantorovich’s OT Problem involving $P$ and $Q$. This implies that the optimal lower coupling $\underline{\alpha}$ is given by $\underline{\alpha}(\text{d}(x,y))=\underline{P}(\text{d}x)\delta_{y=T(x)}$. ∎
@@
A consequence of Theorem 12 is that, for $\epsilon$-contaminated credal sets, LPM and LPK with $\Gamma(\underline{P},\underline{Q})\equiv\Gamma^{\text{geom}}(\underline{P},\underline{Q})$ need not coincide. To see this, notice that Theorem 11 only holds for the restricted LPK (RLPK) in Definition 10. Had we not specified that the joint lower probabilities $\underline{G}\in\Gamma_{R}(\underline{P},\underline{Q})$ are $\epsilon$-contaminations of countably additive joint probabilities $G\in\Delta_{\mathcal{X}\times\mathcal{Y}}$, then Theorem 11 may not have held. Similarly, had we considered generalized Bayes’ conditioning, or other conditioning mechanisms for lower probabilities (Caprio and Seidenfeld, 2023), Theorem 11 may not have held as well.
We conjecture that the lower probability versions of Monge’s and (unrestricted) Kantorovich’s problems not coinciding with one another is a general phenomenon pertaining all types of credal sets, and not only the $\epsilon$-contaminated ones studied here.

### v3

@@ 3.2 Lower Probability Kantorovich’s (LPK) Problem
Let $c:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}_{+}$ be a Borel measurable (cost) function. Given lower probabilities $\underline{P}$ and $\underline{Q}$ on $\mathcal{X}$ and $\mathcal{Y}$, respectively, we want to find the joint lower probability $\underline{\alpha}$ (also called the lower optimal transport plan) on $\mathcal{X}\times\mathcal{Y}$ that solves the following optimization problem
$\operatorname*{arg\,inf}\left\{{\int_{\mathcal{X}\times\mathcal{Y}}c(x,y)\text{d}\underline{\alpha}(x,y):\underline{\alpha}\in\Gamma(\underline{P},\underline{Q})}\right\},$ (12)
@@
Considerations on the choice of a Borel measurable cost function $c$ and of working with a Choquet integral similar to those pointed out below Definition 7 hold also for Definition 9. In Imprecise Probability theory (Augustin et al., 2014; Troffaes and de Cooman, 2014; Walley, 1991) there is not a unique way to perform conditioning (Gong and Meng, 2021; Caprio and Seidenfeld, 2023; Caprio and Gong, 2023), so we need to be extra careful when defining $\Gamma(\underline{P},\underline{Q})$ in (12). In this work, we consider the joint lower probabilities resulting from geometric conditioning, and write $\Gamma(\underline{P},\underline{Q})\equiv\Gamma^{\text{geom}}(\underline{P},\underline{Q})$. In that case, the conditional lower probability resulting from a joint probability $\underline{G}$ is derived as
$\underline{G}(A\mid B)=\frac{\underline{G}(A,B)}{\underline{G}_{\mathcal{Y}}(B)},\quad\forall A\in\mathcal{F},\forall B\text{ s.t. }\underline{G}_{\mathcal{Y}}(B)>0,$ (13)
where $\underline{G}_{\mathcal{Y}}\equiv\underline{Q}$ is the marginal lower probability of $\underline{G}$ on $\mathcal{Y}$, and similarly for $\underline{G}_{\mathcal{X}}\equiv\underline{P}$. The importance of the choice of geometric conditioning is further discussed in Appendix B. Let us mention here that – as pointed out by Gong and Meng (2021) – the agent that chooses the geometric rule as a mechanism to update their belief is a pessimist. In fact, the geometric rule endorses a stringent interpretation of what counts as evidence for both the query ($A$) and conditioning ($B$) events, by admitting only evidence that supports its constituents into the lower conditional probability.
3.3 Restricted Lower Probability Kantorovich’s (RLPK) Problem
If the marginal lower probabilities correspond to the lower envelopes of $\epsilon$-contaminated credal sets, then using joint lower probabilities that can be decomposed as in (13) entails that the elements of $\Gamma(\underline{P},\underline{Q})$ are such that, for all $A\in\mathcal{F}$ and all $B\in\mathcal{G}$ such that $\underline{Q}(B)>0$,
@@ 3.3 Restricted Lower Probability Kantorovich’s (RLPK) Problem
and similarly, $\underline{G}(B\mid A)=\frac{\underline{G}(B,A)}{\underline{P}(A)}=\frac{\underline{G}(A,B)}{(1-\epsilon)P(A)}$.
We can then consider a restricted version of LPK.
@@
Let $c:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}_{+}$ be a Borel measurable (cost) function. Given lower probabilities $\underline{P}$ and $\underline{Q}$ on $\mathcal{X}$ and $\mathcal{Y}$, respectively, we want to find the joint lower probability $\underline{\alpha}$ on $\mathcal{X}\times\mathcal{Y}$ that that solves the following optimization problem
$\operatorname*{arg\,inf}\left\{{\int_{\mathcal{X}\times\mathcal{Y}}c(x,y)\text{d}\underline{\alpha}(x,y):\underline{\alpha}\in\Gamma_{R}(\underline{P},\underline{Q})}\right\},$ (14)
where $\Gamma_{R}(\underline{P},\underline{Q})\subset\Gamma(\underline{P},\underline{Q})$ is the collection of all joint lower probabilities (i) that can be written as an $\epsilon$-contamination of countably additive joint probability measures $G\in\Delta^{\text{ca}}_{\mathcal{X}\times\mathcal{Y}}$, and (ii) whose marginals on $\mathcal{X}$ and $\mathcal{Y}$ are $\underline{P}$ and $\underline{Q}$, respectively.
@@
Notice how working with $\Gamma_{R}(\underline{P},\underline{Q})$ is reminiscent of the covariate shift condition in the Machine Learning literature (Raitoharju, 2022, Section 3.3.4.4). That is, a situation in which there is ambiguity on the marginal distribution of the input features of the model (but not on the conditional distribution of the output, given the input), which may be different (i.e. may have changed) from the one that the model has “seen” during training and validation.
@@
Suppose $\mathcal{X}$ and $\mathcal{Y}$ are separable, so that the elements of $\Delta^{\text{\em ca}}_{\mathcal{X}}$ and $\Delta^{\text{\em ca}}_{\mathcal{Y}}$ are Radon measures. If $\underline{P}$ and $\underline{Q}$ are the lower envelopes of the $\epsilon$-contaminations $\mathcal{P}_{\epsilon}\subseteq\Delta^{\text{\em fa}}_{\mathcal{X}}$ and $\mathcal{Q}_{\epsilon}\subseteq\Delta^{\text{\em fa}}_{\mathcal{Y}}$ of $P\in\Delta^{\text{\em ca}}_{\mathcal{X}}$ and $Q\in\Delta^{\text{\em ca}}_{\mathcal{Y}}$, respectively, then the RLPK of Definition 10 is equivalent to the classical Kantorovich’s OT Problem involving $P$ and $Q$.
@@
$\underline{G}(A,B)$ $=\underline{G}(A\mid B)\underline{Q}(B)=\frac{G(A,B)}{Q(B)}(1-\epsilon)Q(B)$ (15)
$=\underline{G}(B\mid A)\underline{P}(A)=\frac{G(A,B)}{P(A)}(1-\epsilon)P(A)$ (16)
$=(1-\epsilon)G(A,B).$
@@
So we can write $\Gamma_{R}(\underline{P},\underline{Q})=(1-\epsilon)\Gamma(P,Q)=\{(1-\epsilon)G:G\in\Gamma(P,Q)\}$, where set $\Gamma(P,Q)$ is the the collection of all (countably additive) probability measures on $\mathcal{X}\times\mathcal{Y}$ whose marginals on $\mathcal{X}$ and $\mathcal{Y}$ are ${P}$ and ${Q}$, respectively. This shows that $\Gamma_{R}(\underline{P},\underline{Q})$ is nonempty if and only if $\Gamma(P,Q)\neq\emptyset$. In addition, it is easy to see that $\Gamma_{R}(\underline{P},\underline{Q})$ inherits the convexity and the weak-compactness from $\Gamma(P,Q)$. [^9] It is easy to see that $\Gamma(P,Q)$ is convex. In addition, its weak-compactness comes from Prokhorov’s theorem, following our separability assumptions and the well-known fact that $\Gamma(P,Q)$ is weak-closed. In turn,
@@
$\operatorname*{arg\,inf}\left\{{\int_{\mathcal{X}\times\mathcal{Y}}c(x,y)\text{d}\underline{\alpha}(x,y):\underline{\alpha}\in\Gamma_{R}(\underline{P},\underline{Q})}\right\}=$
@@
$\operatorname*{arg\,inf}\left\{{(1-\epsilon)\int_{\mathcal{X}\times\mathcal{Y}}c(x,y)\text{d}{\alpha}(x,y):{\alpha}\in\Gamma(P,Q)}\right\}$ (17)
@@
$=\operatorname*{arg\,inf}\left\{{\int_{\mathcal{X}\times\mathcal{Y}}c(x,y)\text{d}{\alpha}(x,y):{\alpha}\in\Gamma(P,Q)}\right\},$ (18)
@@
where (17) comes from Proposition 4 and our definition of $\Gamma_{R}(\underline{P},\underline{Q})$, and the last equality comes from $c$ being positive. The fact that (18) is the classical Kantorovich’s OT Problem (Kantorovich, 1942) concludes our proof. Notice also that since $\underline{P}$ and $\underline{Q}$ are both countably additive measures normed to $(1-\epsilon)$, the Choquet integrals in this proof coincide with classical Lebesgue-Stieltjes integrals. ∎
Notice that for Theorem 11 too we do not need to implicitly assume that the contaminating parameter $\epsilon$ is the same for both $\mathcal{P}_{\epsilon}$ and $\mathcal{Q}_{\epsilon}$. That is, we could consider $\mathcal{P}_{\epsilon}$ and $\mathcal{Q}_{\epsilon^{\prime}}$, $\epsilon^{\prime}\neq\epsilon$. This because, by (15), we have that $\underline{G}(A,B)=(1-\epsilon^{\prime})G(A,B)$, and, by (16), that $\underline{G}(A,B)=(1-\epsilon)G(A,B)$. But they must be equal to each other, and so $\epsilon=\epsilon^{\prime}$ must hold.
@@
We now give sufficient conditions for the minimizer of (14) to exist, in the context of $\epsilon$-contaminated credal sets. First, we need to introduce the concept of tightness of $\Gamma_{R}(\underline{P},\underline{Q})$.
Definition 12 (Tightness of $\Gamma_{R}(\underline{P},\underline{Q})$).
Let $(\mathcal{X}\times\mathcal{Y},\tau)$ be a Hausdorff space. Let $\Sigma_{\mathcal{X}\times\mathcal{Y}}$ be a $\sigma$-algebra on $\mathcal{X}\times\mathcal{Y}$ that contains $\tau$. That is, every $\tau$-open subset of $\mathcal{X}\times\mathcal{Y}$ is measurable, and $\Sigma_{\mathcal{X}\times\mathcal{Y}}$ is at least as fine as the Borel $\sigma$-algebra on $\mathcal{X}\times\mathcal{Y}$. We say that $\Gamma_{R}(\underline{P},\underline{Q})$ is tight if, for all $\delta\in(0,1]$, there exists a $\tau$-compact set $K_{\delta}\in\Sigma_{\mathcal{X}\times\mathcal{Y}}$ such that, for all $\underline{\alpha}\in\Gamma_{R}(\underline{P},\underline{Q})$, we have that $\underline{\alpha}(K_{\delta})>1-\delta$.
Lemma 13 (Necessary and Sufficient Condition for $\Gamma_{R}(\underline{P},\underline{Q})$ to be Tight).
Let $\mathcal{X},\mathcal{Y}$ be metric spaces, and $\epsilon\in[0,1)$. Then, set $\Gamma_{R}(\underline{P},\underline{Q})$ is tight if and only if set $\Gamma({P},{Q})$ is tight.
Proof.
Suppose $\Gamma_{R}(\underline{P},\underline{Q})$ is tight. Given our assumption that $\mathcal{X}$ and $\mathcal{Y}$ are metric spaces, this implies that $\mathcal{X}$ and $\mathcal{Y}$ are separable. Pick any $\delta\in(0,1]$ and any $\underline{\alpha}\in\Gamma_{R}(\underline{P},\underline{Q})$. Then, by Definition 12, we have that $\underline{\alpha}(K_{\delta})>1-\delta$. By the proof of Theorem 11, we know that there exists $\alpha\in\Gamma({P},{Q})$ such that $\underline{\alpha}(A)=(1-\epsilon){\alpha}(A)$, for all $A\in\Sigma_{\mathcal{X}\times\mathcal{Y}}$. In turn, this implies that $(1-\epsilon){\alpha}(K_{\delta})>1-\delta\iff{\alpha}(K_{\delta})>\frac{1-\delta}{1-\epsilon}$. Now let $\frac{1-\delta}{1-\epsilon}\eqqcolon 1-\gamma$, and put $K_{\delta}\equiv K_{\gamma}$. We obtain ${\alpha}(K_{\gamma})>1-\gamma$. But $\delta$ and $\underline{\alpha}$ were chosen arbitrarily, which allows us to conclude that $\Gamma({P},{Q})$ is tight.
Suppose instead that $\Gamma({P},{Q})$ is tight. As before, given our assumption that $\mathcal{X}$ and $\mathcal{Y}$ are metric spaces, this implies that $\mathcal{X}$ and $\mathcal{Y}$ are separable. Pick any $\delta\in(0,1]$, and any $\alpha\in\Gamma({P},{Q})$. Then, by Definition 12, we have that ${\alpha}(K_{\delta})>1-\delta$. This holds if and only if $(1-\epsilon){\alpha}(K_{\delta})=\underline{\alpha}(K_{\delta})>(1-\epsilon)(1-\delta)$, where $\epsilon\in[0,1)$ is the same parameter of Definition 10. Now let $(1-\epsilon)(1-\delta)\eqqcolon 1-\gamma$, and put $K_{\delta}\equiv K_{\gamma}$. We obtain $\underline{\alpha}(K_{\gamma})>1-\gamma$. But $\delta$ and ${\alpha}$ were chosen arbitrarily, which allows us to conclude that $\Gamma_{R}(\underline{P},\underline{Q})$ is tight. ∎
We are ready for our result.
@@
Let $\mathcal{X},\mathcal{Y}$ be metric spaces, and $\epsilon\in[0,1)$. If $\Gamma_{R}(\underline{P},\underline{Q})$ is tight, and if cost function $c$ in (14) is also lower semicontinuous, then a minimizer for (14) exists.
@@
Let $\Gamma_{R}(\underline{P},\underline{Q})$ be tight. Given our assumption that $\mathcal{X}$ and $\mathcal{Y}$ are metric spaces, this implies that $\mathcal{X}$ and $\mathcal{Y}$ are separable. Ambrosio et al. (2005) show that if $\Gamma(P,Q)$ is tight, and if $c$ is lower semicontinuous, then there is a minimizer for the classical Kantorovich’s OT problem. By Lemma 13, we know that – for any $\epsilon\in[0,1)$ – if $\Gamma_{R}(\underline{P},\underline{Q})$ is tight, then so is $\Gamma(P,Q)$. The proof follows by the equivalence established in Theorem 11. ∎
The tightness condition in Corollary 13.1 is satisfied e.g. when $\mathcal{X}$ and $\mathcal{Y}$ are both Polish spaces. [^10] Separable, completely metrizable topological spaces This is an immediate consequence of Thorpe (2018, Proposition 1.5).
3.4 Equivalence Between LPM And RLPK Problems
@@
Suppose $\mathcal{X}$ and $\mathcal{Y}$ are separable, so that the elements of $\Delta^{\text{\em ca}}_{\mathcal{X}}$ and $\Delta^{\text{\em ca}}_{\mathcal{Y}}$ are Radon measures. Let $\underline{P}$ and $\underline{Q}$ be the lower envelopes of the $\epsilon$-contaminations $\mathcal{P}_{\epsilon}\subseteq\Delta^{\text{\em fa}}_{\mathcal{X}}$ and $\mathcal{Q}_{\epsilon}\subseteq\Delta^{\text{\em fa}}_{\mathcal{Y}}$ of $P\in\Delta^{\text{\em ca}}_{\mathcal{X}}$ and $Q\in\Delta^{\text{\em ca}}_{\mathcal{Y}}$, respectively. When the minimizer $\underline{\alpha}$ of RLPK is such that $\text{d}\underline{\alpha}(x,y)=\underline{P}(\text{\em\text{d}}x)\delta_{y=T(x)}$, then $T$ is an optimal transport map and RLPK is equivalent to LPM.
@@
Given the way we defined $\underline{P}$ and $\underline{Q}$, we have that $\text{d}\underline{\alpha}(x,y)=\underline{P}(\text{d}x)\delta_{y=T(x)}\iff(1-\epsilon)\text{d}{\alpha}(x,y)=(1-\epsilon){P}(\text{d}x)\delta_{y=T(x)}\iff\text{d}{\alpha}(x,y)={P}(\text{d}x)\delta_{y=T(x)}$. Thorpe (2018, Section 1.2) shows that when the minimizer ${\alpha}$ of Kantorovich’s classical OT problem is such that $\text{d}{\alpha}(x,y)={P}(\text{d}x)\delta_{y=T(x)}$, [^11] Conditions sufficient for such a condition can be found in Thorpe (2018, Chapter 4). then $T$ is Monge’s OT map, and Monge’s and Kantorovich’s problems are equivalent. The proof, then, follows by Theorems 8 and 11. ∎
@@
Let $\underline{P}$ and $\underline{Q}$ be the lower envelopes of the $\epsilon$-contaminations $\mathcal{P}_{\epsilon},\mathcal{Q}_{\epsilon}\subseteq\Delta^{\text{\em fa}}_{\mathbb{R}^{d}}$ of $P=\mathcal{N}_{d}(0,\Sigma_{P})$ and $Q=\mathcal{N}_{d}(0,\Sigma_{Q})$, two (countably additive) multivariate Normals on $\mathcal{X}=\mathcal{Y}=\mathbb{R}^{d}$, respectively. Select $c(x,y)=|y-Ax|^{2}/2$, where $A\in\mathbb{R}^{d\times d}$ is invertible. Then, the optimal map that solves LPM is $x\mapsto T(x)$,
$T(x)=(A^{\top})^{-1}\Sigma_{P}^{-1/2}\left(\Sigma_{P}^{1/2}A^{\top}\Sigma_{Q}A\Sigma_{P}^{1/2}\right)^{1/2}\Sigma_{P}^{-1/2}x$ (19)
and the optimal plan that solves RLPK is $\text{d}\underline{\alpha}(x,y)=\underline{P}(\text{\em\text{d}}x)\delta_{y=T(x)}$.
@@
Notice that $\mathbb{R}^{d}$ is separable. Galichon (2016) shows that if $P=\mathcal{N}_{d}(0,\Sigma_{P})$, $Q=\mathcal{N}_{d}(0,\Sigma_{Q})$, and $c(x,y)=|y-Ax|^{2}/2$, then Monge’s (classical) OT map is the one in (19), and also that $\text{d}{\alpha}(x,y)={P}(\text{d}x)\delta_{y=T(x)}$, so that Monge’s and Kantorovich’s (classical) problems are equivalent.
Now, by Theorem 8, we have that if $\underline{P}$ and $\underline{Q}$ are lower envelopes of the $\epsilon$-contaminations $\mathcal{P}_{\epsilon},\mathcal{Q}_{\epsilon}\subseteq\Delta^{\text{fa}}_{\mathbb{R}^{d}}$ of $P=\mathcal{N}_{d}(0,\Sigma_{P})$ and $Q=\mathcal{N}_{d}(0,\Sigma_{Q})$, respectively, then LPM coincides with the classical Monge’s OT Problem involving $P$ and $Q$. In turn, this implies that the map in (19) is also the OT map between $\underline{P}$ and $\underline{Q}$. In addition, by Theorem 11, we know that RLPK coincides with the classical Kantorovich’s OT Problem involving $P$ and $Q$. This implies that the optimal lower coupling $\underline{\alpha}$ is given by $\text{d}\underline{\alpha}(x,y)=\underline{P}(\text{d}x)\delta_{y=T(x)}$. ∎
@@
A consequence of Theorem 14 is that, for $\epsilon$-contaminated credal sets, LPM and LPK with $\Gamma(\underline{P},\underline{Q})\equiv\Gamma^{\text{geom}}(\underline{P},\underline{Q})$ need not coincide. To see this, notice that Theorem 11 only holds for the restricted LPK (RLPK) in Definition 10. Had we not specified that the joint lower probabilities $\underline{G}\in\Gamma_{R}(\underline{P},\underline{Q})$ are $\epsilon$-contaminations of countably additive joint probabilities $G\in\Delta^{\text{ca}}_{\mathcal{X}\times\mathcal{Y}}$, then Theorem 11 may not have held. Similarly, had we considered generalized Bayes’ conditioning, or other conditioning mechanisms for lower probabilities (Caprio and Seidenfeld, 2023), Theorem 11 may not have held as well.
Whether this is a phenomenon pertaining only to $\epsilon$-contaminated credal sets, or a more general one, will be the subject of future studies.